Hayley Company just paid a dividend of $2 today, and is expected to pay a dividend in year 1 of $2.3, a dividend in year 2 of $2.4, a dividend in year 3 of $3.5, a dividend in year 4 of $3.9, and a dividend in year 5 of $4.8. After year 5, dividends are expected to grow at the rate of 0.05 per year. An appropriate required return for the stock is 0.12. Using the different-stage growth model, the stock should be worth __________ today.

*Round your answer to TWO decimal places.
Answer #1

D1 = $2.3

D2 = $2.4

D3= $3.5

D4 = $3.9

D5 = $4.8

D6 = $5.04

P5 = D6/ Re - G

= 5.04/0.12 - 0.05

= $72

The stock should be worth :

$2.3/1.12 + 2.4 /1.12^2 + 3.5/1.12^3 + 3.9/1.12^4 + 4.8/1.12^5 + 72/1.12^5

= 2.0536 + 1.9133 + 2.4912 + 2.4785 + 2.7236 + $40.8547

= $52.5149

=$52.51 (ROUNDED OFF TO TWO DECIMAL PLACES)

The stock should be worth $52.51 today.